Student Council wants to rent a banquet hall for the prom. Caesar's charges $800 to rent the room and $30 for each meal. Venus Hall charges $500 to rent the room and $35 for each meal. How many guests must attend the prom for the costs of the two halls to be the same?
Answer by checkley77(12569)

800+30X=500+35X
30X-35X=500-800
-5X=-300
X=-300/-5
X=60 GUESTS ARE REQUIRED FOR THE COSTS TO BE EQUAL.
PROOF:
800+30*60=500+35*60
800+1800=500+2100
2600=2600